OpenStudy (anonymous):

Which is the best equation to use to solve this problem? The perimeter of a rectangle is 90 ft. The ratio of the length to the width is 8: 7. What are the dimensions of the rectangle? A. 8x+7x=90 B. 8x+8x+7x+7x=90 C. 8x/7x=90 D. 2x+2x=90

OpenStudy (unklerhaukus):

2(a+b)=90 8a=7b solve the system

OpenStudy (anonymous):

i dont understand

OpenStudy (mathmale):

UR has demonstrated how this word problem can be reduced to two equations in a and b, with a and b representing the length and width of the rectangle, respectively. It may be easier for you to grasp what's happening here if you'd draw this rectangle. Label the longer side " a " and the shorter side " b " (noting that you could just as well have used x and y or some other literals). What does perimeter mean? What does it mean when we say that the ratio of length to width is 8/7?

OpenStudy (mathmale):

Write the formula for perimeter in terms of a and b or x and y, whichever you prefer. Normally the length of a rectangle is the longer side. We could then literally write an equation showing that the ratio of length to width is 8:7: \[\frac{ length }{ width }=\frac{ 8 }{ 7 }\]. You need to find a relationship between a and b in the form of a linear equation (Uncle R has already done that for you).

OpenStudy (mathmale):

PP: Hope you'll respond in some way.

OpenStudy (unklerhaukus):

2(a+8a/7)=90

OpenStudy (anonymous):

I am still confused but i am going to guess that it is A.

OpenStudy (unklerhaukus):

that cant be right , its not the formula for the perimeter of a rectangle

OpenStudy (mathmale):

PP: You're free to do as you wish, but I did suggest steps for you to take. Please refer back to my first two postings.

OpenStudy (mathmale):

I'd suggest y ou ask questions, even rudimentary ones, to help yourself get started.

OpenStudy (anonymous):

my second guess would be C.

OpenStudy (mathmale):

PP: I really want to be of help, but choose not to respond to "guesses." I need to know and understand the reasoning you're using to arrive at answers.

OpenStudy (anonymous):

Well i know a ratio is a fraction and because 8:7 is a ratio it is a fraction. 8/7. so that is why i am guessing C

OpenStudy (mathmale):

PP: As Uncle R pointed out earlier, we have two "simultaneous equations" and must use both in finding a solution (that is, finding the values of a and b). have you solved simult eqns before? if so, which method would you apply here?

OpenStudy (mathmale):

Here are your "simultaneous equations" (or "system of equations"). Your job is to find either a or b first. 2(a+b)=90 8a=7b Note that the second can be "solved for a," and that a=(7b)/8 . Substitute (7b)/8 for a in the other equation. Type out the resulting equation, please.

OpenStudy (mathmale):

PP: Please respond so that I'll know you're still involved in finding a solution to this problem.

OpenStudy (anonymous):

Sorry my computer is slow

OpenStudy (anonymous):

The answer is B